An archer, about to shoot an arrow, is applying a force of +330 N to a drawn bowstring. The bow behaves like an ideal spring whose spring constant is 510 N/m. What is the displacement of the bowstring?
______________m
Answer #1
f=kx

x is the displacement

x=330/510 =0.647 m = 11/17 m